# On finite group scheme-theoretical categories, II

### v1

@@ 3.3. The first equivalence ${\rm Corep}_{k}(A,\xi^{-1})\cong\mathscr{A}$
(3.22) ${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:={\rm p}{}^{(A,\xi^{-1})}_{*}{\rm t}^{*},\,\,\,{\rm and}\,\,\,{\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:=\rm t_{*}^{((B,\beta),(C,\gamma))}{\rm p}^{*}.$
@@
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{\cong}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\cong}{\rm Corep}_{k}(A,\xi^{-1}),$
@@
Set $Q:=Q_{\xi}$, so $Q={\rm S}\left(Q_{\xi^{-1}}^{-1}\right)$ (2.9). Recall (2.31) that for any $(V,\ell)\in{\rm Corep}_{k}(A,\xi^{-1})$, we have $(V,\widetilde{\ell})\in{\rm Corep}(A,\xi)_{k}$, where
~$\widetilde{\ell}:V\to V\otimes\mathscr{O}(A)_{\xi},\,\,\,v\mapsto v^{0}\otimes{\rm S}\left(v^{-1}\right)Q.$
For any $(V,\ell)\in{\rm Corep}_{k}(A,\xi^{-1})$, consider the subspace
@@
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{\cong}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\cong}{\rm Corep}_{k}(A,\xi^{-1}),$
@@
3.4. The second equivalence ${\rm Corep}_{k}(A,\xi^{-1})\cong\mathscr{A}$
@@
Also, for any $V\in{\rm Corep}_{k}(A,\xi^{-1})$, consider the $k$-linear isomorphisms
@@
${\rm F}:{\rm Corep}_{k}(A,\xi^{-1})\xrightarrow{\cong}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@ 4.2. Biequivariant sheaves on $(G,\omega)$
(4.7) ${\rm w}_{g}:=\omega_{1}^{g^{-1}}\cdot\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)\cdot\omega_{3}\in C^{2}(G,\mathbb{G}_{m}).$
Also, let $p_{1}:H\times K\twoheadrightarrow H$ and $p_{2}:H\times K\twoheadrightarrow K$ be the projection morphisms, and define the $2$-cochain
@@
    ${\rm Ind}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
@@
    ${\rm F}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
@@
    $\mathfrak{j}_{g*}:{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g})\xrightarrow{\cong}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(Z,\omega_{Z})$.
@@ 5.1. The simple objects of $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
    $\mathbf{Ind}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\mathscr{M}}_{Z},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
@@
Thus, there exists $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$. Assume that $\mathscr{I}(Z^{\prime})S=0$ for some $Z^{\prime}\in Y(k)$, $Z^{\prime}\neq Z$. Then $\mathscr{O}(G)S=0$ (since $\mathscr{O}(G)=\mathscr{I}(Z)+\mathscr{I}(Z^{\prime})$), a contradiction. It follows that there exists a unique $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$, i.e., $S\in{\mathscr{M}}_{Z}$, so the claim follows from (1).
@@ 5.2. Projectives in $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
$1\to Z^{\circ}\xrightarrow{i_{Z^{\circ}}}Z\mathrel{\mathop{\rightleftarrows}^{\pi_{Z}}_{q_{Z}}}Z(k)\to 1$
@@
    For any simple $V\in{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})$, we have an ${\mathscr{M}}$-isomorphism
@@
On the other hand, note that for any simple $W\in{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})$, the objects $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ and $\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ have the same composition factors as $\mathscr{O}(G)$-modules. Hence, if $V$ and $W$ are nonisomorphic simples in ${\rm Corep}_{k}(L^{g},\xi_{g}^{-1})$, then
@@
    $\mathbf{F}_{Z}:{\rm Corep}_{k}(L^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\mathscr{M}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right),$
    where, we have (see (3.27)-(3.28))
    $\lambda^{g}_{V}:=\left({\rm id}\otimes\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}\right)\lambda_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right),\,\,\rho^{g}_{V}:=\left(\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}^{\otimes 2}\right)\rho_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right).$
@@
Take $K=1$. Then $Y=H\backslash G$, and for any $Z\in Y(k)$ with representative $g\in G(k)$, we have $Z=Hg$, $L^{g}=1$, and
$\mathbf{F}_{Hg}:{\rm Vec}\xrightarrow{\cong}{\rm Coh}^{\left(\left(H,\psi\right),\left(1,1\right)\right)}(Hg),\,\,\,V\mapsto\mathscr{O}(Hg)\otimes_{k}V,$
@@ 5.3. Fiber functors on ${\mathscr{C}}(G,\omega,H,\psi)$
Let ${\mathscr{M}}:=\mathscr{M}\left(\left(H,\psi\right),\left(K,\eta\right)\right)$ be an indecomposable exact module category over ${\mathscr{C}}$. By Theorem 5.1, ${\mathscr{M}}\cong{\rm Vec}$ if and only if ${\mathscr{M}}={\mathscr{M}}_{HK}$ and ${\rm Corep}_{k}(H\cap K,\xi^{-1})\cong{\rm Vec}$. Thus, the statement follows from the fact that ${\mathscr{M}}={\mathscr{M}}_{HK}$ if and only if $\iota_{HK*}$ is an equivalence, i.e., if and only if $G=HK$. ∎
@@ 6. The structure of ${\mathscr{C}}\left(G,\omega,H,\psi\right)$
Fix a group scheme-theoretical category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$2.11. For any closed point $g\in G(k)$, let $H^{g}:=H\cap gHg^{-1}$. Note that $\xi_{1}=1$.
@@
    $\mathbf{Ind}_{Z}:{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})\xrightarrow{\cong}\mathscr{C}_{Z},$
@@
    $\mathbf{Ind}_{H}:{\rm Corep}_{k}(H)\xrightarrow{\cong}\mathscr{C}_{H},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})$ is simple, and simple objects of ${\mathscr{C}}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$. Moreover, we have a direct sum decomposition of abelian categories
@@
    For any $V\in{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})$, we have $\mathbf{Ind}_{Z}(V)^{*}\cong\mathbf{Ind}_{Z^{-1}}(V^{*})$, where $Z^{-1}\in Y(k)$ such that $g^{-1}\in Z^{-1}(k)$.
@@
    $\mathbf{F}_{Z}:{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})\xrightarrow{\cong}{\mathscr{C}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right).$
@@
    $\mathbf{F}_{H}:{\rm Corep}_{k}(H)\xrightarrow{\cong}\mathscr{C}_{H},\,\,\,V\mapsto\iota_{H*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right).$
@@
    For any $V\in{\rm Corep}_{k}(H^{g},\xi_{g}^{-1})$, we have $\mathbf{F}_{Z}(V)^{*}\cong\mathbf{F}_{Z^{-1}}(V^{*})$.
@@
$|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}\left(P(\mathbf{1})\right)$
$\cong$ $\mathscr{O}(G^{\circ})\otimes\mathbf{F}\left(P(\mathbf{1})^{*}\right)\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}\left(P(\mathbf{1})\right)^{*}$
$\cong$ $\mathscr{O}(G^{\circ})^{*}\otimes\mathbf{F}\left(P(\mathbf{1})\right)^{*}\cong\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}\left(P(\mathbf{1})\right)\right)^{*}$
$\cong$ $|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})^{*},$

### v2

@@ 3.3. The first equivalence ${\rm Rep}(A,\xi^{-1})_{k}\simeq\mathscr{A}$
(3.22) ${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:={\rm p}{}^{(A,\xi^{-1})}_{*}{\rm t}^{*}\,\,\,{\rm and}\,\,\,{\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:=\rm t_{*}^{((B,\beta),(C,\gamma))}{\rm p}^{*}.$
@@
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},$
@@
Set $Q:=Q_{\xi}$, so $Q={\rm S}\left(Q_{\xi^{-1}}^{-1}\right)$ (2.10). Recall (2.33) that for any $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, we have $(V,\widetilde{\ell})\in{\rm Rep}_{k}(A,\xi)$, where
$\widetilde{\ell}:V\to V\otimes\mathscr{O}(A)_{\xi},\,\,\,v\mapsto v^{0}\otimes{\rm S}\left(v^{-1}\right)Q.$
For any $(V,\ell)\in{\rm Rep}(A,\xi^{-1})_{k}$, consider the subspace
@@
${\rm Ind}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@
${\rm Res}_{(A,\xi^{-1})}^{((B,\beta),(C,\gamma))}:{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right)\xrightarrow{\simeq}{\rm Rep}(A,\xi^{-1})_{k},$
@@
3.4. The second equivalence ${\rm Rep}(A,\xi^{-1})_{k}\simeq\mathscr{A}$
@@
Also, for any $V\in{\rm Rep}(A,\xi^{-1})_{k}$, consider the $k$-linear isomorphisms
@@
${\rm F}:{\rm Rep}(A,\xi^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{((B,\beta),(C,\gamma))}\left(A\backslash(B\times C),\Phi\right),$
@@ 4.2. Biequivariant sheaves on $(G,\omega)$
(4.7) ${\rm w}_{g}:=\omega_{1}^{g^{-1}}\cdot\left({\rm id}\otimes{\rm Ad}g^{-1}\right)\left(\omega_{2}^{-1}\right)\cdot\omega_{3}\in C^{2}(G,\mathbb{G}_{m})$
(see (2.13)). Also, let $p_{1}:H\times K\twoheadrightarrow H$ and $p_{2}:H\times K\twoheadrightarrow K$ be the obvious projection morphisms, and define the $2$-cochain
@@
    ${\rm Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
@@
    ${\rm F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g}),$
@@
    $\mathfrak{j}_{g*}:{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(L^{g}\backslash(H\times K),\Phi_{g})\xrightarrow{\simeq}{\rm Coh}^{\left((H,\psi),(K,\eta)\right)}(Z,\omega_{Z})$.
@@ 5.1. The simple objects of $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
    $\mathbf{Ind}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{M}$, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$.
@@
Thus, there exists $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$. Assume that $\mathscr{I}(Z^{\prime})S=0$ for some $^{\prime}\in Y(k), ZZ$$^{\prime}\neq Z$. Then $\mathscr{O}(G)S=0$ (since $\mathscr{O}(G)=\mathscr{I}(Z)+\mathscr{I}(Z^{\prime})$), a contradiction. It follows that there exists a unique $Z\in Y(k)$ such that $\mathscr{I}(Z)S=0$, i.e., $S\in{\mathscr{M}}_{Z}$, so the claim follows from (1).
@@ 5.2. Projectives in $\mathscr{M}\left((H,\psi),(K,\eta)\right)$
$1\to Z^{\circ}\xrightarrow{i_{Z^{\circ}}}Z\mathrel{\mathop{\rightleftarrows}^{\pi_{Z}}_{q_{Z}}}Z(k)\to 1.$
@@
    For any simple $V\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, we have an $\mathscr{M}$}-isomorphism
@@
On the other hand, note that for any simple $W\in{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, the objects $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ and $\mathbf{F}_{Z}\left(P_{(L^{g},\xi_{g}^{-1})}(W)\right)$ have the same composition factors as $\mathscr{O}(G)$-modules. Hence, if $V$ and $W$ are nonisomorphic simples in ${\rm Rep}(L^{g},\xi_{g}^{-1})_{k}$, then
@@
    $\mathbf{F}_{Z}:{\rm Rep}(L^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{M}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right),$
    where
    $\lambda^{g}_{V}:=\left({\rm id}\otimes\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}\right)\lambda_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right),\,\,\rho^{g}_{V}:=\left(\left(\mathfrak{j}_{g}^{\sharp}\right)^{-1}\otimes{\rm id}^{\otimes 2}\right)\rho_{V}\left(\mathfrak{j}_{g}^{\sharp}\otimes{\rm id}\right)$
    (see (3.27)-(3.28)).
@@
[…]
$\mathbf{F}_{Hg}:{\rm Vec}\xrightarrow{\simeq}{\rm Coh}^{\left(\left(H,\psi\right),\left(1,1\right)\right)}(Hg),\,\,\,V\mapsto\mathscr{O}(Hg)\otimes_{k}V,$
@@ 5.3. Fiber functors on ${\mathscr{C}}(G,\omega,H,\psi)$
Let ${\mathscr{M}}:=\mathscr{M}\left(\left(H,\psi\right),\left(K,\eta\right)\right)$ be an indecomposable exact module category over $\mathscr{C}$}. By Theorem 5.1, ${\mathscr{M}}\simeq{\rm Vec}$ if and only if ${\mathscr{M}}={\mathscr{M}}_{HK}$ and ${\rm Corep}_{k}(\mathscr{O}(H\cap K)_{\xi^{-1}})\simeq{\rm Vec}$. Thus, the statement follows from the fact that ${\mathscr{M}}={\mathscr{M}}_{HK}$ if and only if $\iota_{HK*}$ is an equivalence, i.e., if and only if $G=HK$. ∎
@@ 6. The structure of ${\mathscr{C}}\left(G,\omega,H,\psi\right)$
Fix a group scheme-theoretical category ${\mathscr{C}}:={\mathscr{C}}\left(G,\omega,H,\psi\right)$ §2.12. For any closed point $g\in G(k)$, let $H^{g}:=H\cap gHg^{-1}$. Note that $\xi_{1}=1$.
Recall that $Y=G/(H\times K)$4.1).
@@
    $\mathbf{Ind}_{Z}:{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}\mathscr{C}_{Z},$
@@
    $\mathbf{Ind}_{H}:{\rm Rep}(H)_{k}\xrightarrow{\simeq}\mathscr{C}_{H},$
@@
    There is a bijection between equivalence classes of pairs $(Z,V)$, where $Z\in Y(k)$ is a closed point with representative $g\in Z(k)$, and $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$ is simple, and simple objects of $\mathscr{C}$}, assigning $(Z,V)$ to $\mathbf{Ind}_{Z}(V)$. Moreover, we have a direct sum decomposition of abelian categories
@@
    For any $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have $\mathbf{Ind}_{Z}(V)^{*}\cong\mathbf{Ind}_{Z^{-1}}(V^{*})$, where $Z^{-1}\in Y(k)$ such that $g^{-1}\in Z^{-1}(k)$.
@@
    $\mathbf{F}_{Z}:{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}\xrightarrow{\simeq}{\mathscr{C}}_{Z},\,\,\,V\mapsto\iota_{Z*}\left(\mathscr{O}(Z)\otimes_{k}V,\lambda^{g}_{V},\rho^{g}_{V}\right).$
@@
    $\mathbf{F}_{H}:{\rm Rep}(H)_{k}\xrightarrow{\simeq}\mathscr{C}_{H},\,\,\,V\mapsto\iota_{H*}\left(\mathscr{O}(H)\otimes_{k}V,\lambda^{1}_{V},\rho^{1}_{V}\right).$
@@
    For any $V\in{\rm Rep}(H^{g},\xi_{g}^{-1})_{k}$, we have $\mathbf{F}_{Z}(V)^{*}\cong\mathbf{F}_{Z^{-1}}(V^{*})$.
@@
$|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)$
$\cong$ $\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})^{*}\right)\cong\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)^{*}$
$\cong$ $\left(\mathscr{O}(G^{\circ})\otimes\mathbf{F}_{H}\left(P(\mathbf{1})\right)\right)^{*}\cong|H^{\circ}|P_{{\mathscr{C}}}(\mathbf{1})^{*},$